OpenStudy (anonymous):

How many integer values are there for k for which 4x2 + kxy – 9y2 is factorable? HELP :)

OpenStudy (raden):

look at the expanding form of the polynomials bellow : (4x + 3y)(x - 3y) = 4x^2 - 9xy - 9y^2 (x + 3y)(4x - 3y) = 4x^2 + 9xy - 9y^2 (2x + 3y)(2x - 3y) = 4x^2 - 9y^2 now, can u determine the values all k's are there which satisfying is factorable ?